OpenStudy (anonymous):

\[f(x)=-\sqrt{3-x}\]

OpenStudy (anonymous):

need help finding the domain

OpenStudy (anonymous):

solve \(3-x\geq 0\) in one step

OpenStudy (anonymous):

\[x \ge-3\]

OpenStudy (anonymous):

\[x \le-3\]